# Approximation and bounding techniques for the Fisher-Rao distances

### v2

@@ Abstract
The Fisher-Rao distance between two probability distributions of a finite-dimensional parametric statistical model is defined as the Riemannian geodesic distance induced by the Fisher information metric. The Fisher-Rao distance is guaranteed by construction to be invariant under diffeomorphisms of both the sample space and the parameter space of the statistical model. In order to calculate the Fisher-Rao distance in closed-form, we need (1) to elicit a formula for the Fisher-Rao geodesics, and (2) to integrate the Fisher length element along those geodesics. These calculations turn out to be difficult tasks for some statistical models like the multivariate elliptical distributions which include the ubiquitous multivariate normal distributions. In this work, we consider several numerically robust approximation and bounding techniques for the Fisher-Rao distances: First, we report generic upper bounds on Fisher-Rao distances based on closed-form 1D Fisher-Rao distances of submodels. Second, we describe several generic approximation schemes depending on whether the Fisher-Rao geodesics or pregeodesics are available in closed-form or not. In particular, we obtain a generic method to guarantee an arbitrarily small additive error on the approximation provided that Fisher-Rao pregeodesics and tight lower and upper bounds are available. This case applies to the multivariate normal distributions. Third, we consider the case of Fisher metrics being Hessian metrics, and report generic tight upper bounds on the Fisher-Rao distances by square roots of Jeffreys-Bregman divergences which correspond to the Fisher-Rao energies of dual curves in information geometry. Uniparametric and biparametric statistical models always have Fisher Hessian metrics, and in general a simple test allows to check whether the Fisher information matrix yields a Hessian Fisher metric or not. Fourth, we consider elliptical distribution families and show how to apply the above techniques to these models. We also propose two new distances based either on the Fisher-Rao lengths of curves serving as proxies of Fisher-Rao geodesics, or based on the Birkhoff/Hilbert projective cone distance. Last, we consider an alternative group-theoretic approach for statistical transformation models based on the notion of maximal invariant which yields insights on the structures of the Fisher-Rao distance formula and its degrees of freedom which may be used fruitfully in applications.
@@
Keywords: Fisher information; Riemannian geometry; Malahanobis distance; (pre)geodesics; Rao’s distance; information geometry; Hessian metric/manifold; Bregman divergence; isometric embedding; elliptical distribution; Birkhoff/Hilbert projective cone geometry; transformation model; maximal invariant.
@@ 1 Introduction
In §5, we consider Fisher metrics which are Hessian metrics [96], and report upper bounds (Proposition 8) which are square roots of Jeffreys-Bregman divergences on the Fisher-Rao distances and correspond to Fisher-Rao energies of dual curves well-studied in information geometry (Proposition 7). Those square roots of Jeffreys-Bregman upper bounds are tight at infinitesimal scales. Examples of Hessian Fisher metrics are Fisher metrics obtained from exponential or mixture families, 2D elliptical distribution families, etc. In practice, a simple test on the Fisher information matrix (Property 1) allows to check whether the corresponding Fisher metric is a Hessian metric or not.
@@ 2.1 Preliminary: Riemannian geometry of domains
$IVC:\quad\left\{\begin{array}[]{l}\ddot{\gamma}_{k}+\sum_{i,j}\Gamma_{ij}^{k}\,\dot{\gamma}_{i}\dot{\gamma}_{j}=0\\ \gamma(0),\dot{\gamma}(0)\in T_{\gamma(0)}\end{array}\right.,$
@@
$BVC:\quad\left\{\begin{array}[]{l}\ddot{\gamma}_{k}+\sum_{i,j}\Gamma_{ij}^{k}\,\dot{\gamma}_{i}\dot{\gamma}_{j}=0\\ \gamma(0),\gamma(1)\end{array}\right..$
@@ 2.2 Definition of the Fisher-Rao distance and Fisher-Rao geodesics
Those Fisher information matrices $I(\lambda)$, $I(\lambda^{\prime})$ and $I(\lambda^{\prime\prime})$ define the same underlying Fisher Riemannian metric $g^{\mathrm{Fisher}}$ and length element.
@@
Stigler [100] noticed that Hotelling historically first considered the Fisher-Rao distance in a handwritten note [54] dated in 1929.
@@ 3.2 Fisher-Rao distances for products of stochastically independent models
For $m$ independent statistical models $\mathcal{M}_{1}=\{p_{\theta^{1}}(x_{1})\}$ , normal-$\ldots$, $\mathcal{M}_{m}=\{p_{\theta^{m}}(x_{m})\}$ with corresponding variables $x_{i}\in\mathbb{R}^{d_{i}}$ and parameters $\theta^{i}\in\mathbb{R}^{m_{i}}$ (with $d=\sum_{i}d_{i}$ and $m=\sum_{i}m_{i}$), and Fisher-Rao’s distances $\rho_{\mathcal{M}_{1}}$, normal-$\ldots$, $\rho_{\mathcal{M}_{m}}$, we have the Fisher-Rao distance on the product model =$\mathcal{M}=\{p_{\theta}(x)\propto\prod_{i=1}^{m}p_{\theta^{i}}(x_{i}){\ :\ }\theta\in\mathbb{R}^{m}\}$ which is given by
@@ 3.3 The Fisher-Manhattan upper bound for the Fisher-Rao distances
Figure 2: Illustration of the Fisher-Manhattan upper bound for a bidimensional statistical model: We calculate the 1D Fisher-Rao distances along the edges and then find the shortest path on the weighted graph between node $\theta_{0}$ and node $\theta_{1}$.
@@
Figure 3: Illustrating the Fisher-Manhattan upper bound for the Fisher-Rao manifold of univariate normal distributions ($m=2$): The Fisher-Rao geodesic between $N(\mu_{0},\sigma_{0})$ and $N(\mu_{1},\sigma_{1})$ is shown in red. The Fisher-Manhattan upper bound is derived from a square $((\mu_{0},\sigma_{0}),(\mu_{0},\sigma_{1}),(\mu_{1},\sigma_{1}),(\mu_{1},\sigma_{0}))$. The Fisher-Manhattan upper bound is $\min(\rho((\mu_{0},\sigma_{0}),(\mu_{0},\sigma_{1}))+\rho((\mu_{0},\sigma_{1}),(\mu_{1},\sigma_{1})),(\mu_{0},\sigma_{0}),(\mu_{1},\sigma_{0})+\rho((\mu_{1},\sigma_{0}),(\mu_{1},\sigma_{1})))$, i.e., the minimum length of the two green and blue path lengths. Tissot indicatrices of the Fisher metric are displayed at 2D grid locations.
@@ 5 Tight upper bounds on Fisher-Rao distances: The case of Hessian metrics
When the Fisher metric $g^{\mathrm{Fisher}}$ of a statistical model =$\mathcal{M}=\{p_{\theta}{\ :\ }\theta\in\Theta\}$ is a Hessian metric, the Fisher-Rao manifold $(\mathcal{M},g^{\mathrm{Fisher}})$ is called a Hessian manifold [96]. Dually flat spaces in information geometry [10, 4] are special Hessian manifolds which admit global coordinate charts and dual potential functions. Recall that the Fisher metric is defined in the global chart by the Fisher information matrix $I(\theta)=-E_{p_{\theta}}\left[\nabla^{2}_{\theta}\log p_{\theta}(x)\right]$. When the Fisher metric is Hessian, we have $G(\theta)=I(\theta)=\nabla^{2}F(\theta)$ for a strictly convex and smooth potential function $F(\theta)$. A dually flat space admits as a canonical divergence a Bregman divergence $B_{F}(\theta_{1}:\theta_{2})=F(\theta_{1})-F(\theta_{2})-(\theta_{1}-\theta_{2})^{\top}\nabla F(\theta_{2})$ induced by a potential Legendre-type convex function $F(\theta)$ (defined up to a modulo affine terms), and two torsion-free flat connections $\nabla$ and $\nabla^{*}$ such that +*$\frac{\nabla+\nabla^{*}}{2}$ coincides with the Levi-Civita connection. These dual connections $\nabla$ and $\nabla^{*}$ coupled to the Hessian Fisher metric $g^{\mathrm{Fisher}}$ allows to connect any two points $p_{\theta_{1}}$ and $p_{\theta_{2}}$ on the manifold either by $\nabla$-geodesics $\gamma_{p_{\theta_{1}}p_{\theta_{2}}}$ or by $\nabla^{*}$-geodesics $\gamma_{p_{\theta_{1}}p_{\theta_{2}}}^{*}$ (Figure 6).
@@
Many statistical models yield Fisher metrics: For example, exponential families [14], mixture families [4], all statistical models of order $m=2$ [72, 24, 7], and some elliptical distribution families [72, 73], etc. In general, there are obstructions for a $m$-dimensional Riemannian metric $g$ to be Hessian [5]. Thus it is remarkable that exponential and mixture families of arbitrary order yield Hessian Fisher metrics. One can test whether a given metric $g$ given in a coordinate system $\theta$ by $G(\theta)=[g_{ij}(\theta)]$ is Hessian or not as follows:
Property 1 (Hessian metric test [96])
A $m$-dimensional Riemannian metric $g$ is Hessian if and only if:
@@
This property holds for any parameterization $\theta$ of the metric $g$.
@@
Property 2 (1d Fisher metrics are Hessians)
@@
A 2D Hessian metric $g$ can always be expressed by a diagonal matrix in some coordinate system by a change of parameterization: Indeed, let $\theta=(\theta_{1},\theta_{2})$ and $\eta=(\eta_{1},\eta_{2})$ be the dual affine normal-$\nabla$- and $\nabla^{*}$-coordinate systems of the flat connections normal-$\nabla$ and $\nabla^{*}$. The Fisher metric is expressed in $\theta$ as $g(\theta)=\nabla^{2}F(\theta)$ and in $\eta$ as $g(\eta)=\nabla^{2}F^{*}(\eta)$. Then we can consider the mixed parameterizations $\xi=(\theta_{1},\eta_{2})$ or $\xi^{\prime}=(\eta_{1},\theta_{2})$ and obtain diagonal matrices for the Fisher information matrices with respect to those mixed parameterizations [4, 75].

### v3

@@ Abstract
The Fisher-Rao distance between two probability distributions of a finite-dimensional parametric statistical model is defined as the Riemannian geodesic distance induced by the Fisher information metric. The Fisher-Rao distance is guaranteed by construction to be invariant under diffeomorphisms of both the sample space and the parameter space of the statistical model. In order to calculate the Fisher-Rao distance in closed-form, one proceeds by (1) eliciting a formula for the Fisher-Rao geodesics, and (2) integrating the Fisher-Rao length element along those geodesics. These calculations turn out to be difficult tasks for some statistical models like the multivariate elliptical distributions which include the class of multivariate normal distributions. In this work, we consider several numerically robust approximation and bounding techniques for the Fisher-Rao distances: First, we report generic upper bounds on the Fisher-Rao distances based on closed-form Fisher-Rao distances of univariate submodels. Second, we describe several generic approximation schemes depending on whether the Fisher-Rao geodesics or pregeodesics are available in closed-form or not. In particular, we obtain a generic method to guarantee an arbitrarily small additive error on the approximation provided that Fisher-Rao pregeodesics and tight lower and upper bounds are available. This case applies to the class of multivariate normal distributions. Third, we consider the case of Fisher metrics being Hessian metrics, and report generic tight upper bounds on the Fisher-Rao distances as square roots of Jeffreys-Bregman divergences which correspond to the Fisher-Rao energies of dual curves in information geometry. Uniparametric and biparametric statistical models always have Fisher Hessian metrics. Fourth, we consider elliptical distribution families and show how to apply the above approximation and bounding techniques to these models. We also propose two new distances based either on the Fisher-Rao lengths of curves serving as proxies of Fisher-Rao geodesics, or based on the Birkhoff/Hilbert projective cone distance. Finally, we consider an alternative group-theoretic approach for statistical transformation models based on the notion of maximal invariant which yields insights on the structures of the Fisher-Rao distance formula and its degrees of freedom which may be used fruitfully in applications.
@@
Keywords: Fisher information; Riemannian geometry; Malahanobis distance; geodesics; Rao’s distance; information geometry; Hessian manifold; Bregman divergence; isometric embedding; elliptical distribution; Birkhoff/Hilbert projective cone geometry; transformation model; maximal invariant.
@@ 1 Introduction
In §5, we consider Fisher metrics which are Hessian metrics [108], and report upper bounds (Proposition 8) which are square roots of Jeffreys-Bregman divergences on the Fisher-Rao distances and correspond to Fisher-Rao energies of dual curves well-studied in information geometry (Proposition 7). Those square roots of Jeffreys-Bregman upper bounds are tight at infinitesimal scales. Examples of Hessian Fisher metrics are Fisher metrics obtained from exponential or mixture families, 2D elliptical distribution families, etc.
@@ 2.1 Preliminary: Riemannian geometry of domains
$\mathrm{IVC}:\quad\left\{\begin{array}[]{l}\ddot{\gamma}_{k}+\sum_{i,j}\Gamma_{ij}^{k}\,\dot{\gamma}_{i}\dot{\gamma}_{j}=0\\ \gamma(0),\dot{\gamma}(0)\in T_{\gamma(0)}\end{array}\right.,$
@@
$\mathrm{BVC}:\quad\left\{\begin{array}[]{l}\ddot{\gamma}_{k}+\sum_{i,j}\Gamma_{ij}^{k}\,\dot{\gamma}_{i}\dot{\gamma}_{j}=0\\ \gamma(0),\gamma(1)\end{array}\right..$
@@
For two positive definite matrices $P$ and $Q$, let $P\succeq Q$ denote the Löwner partial ordering. That is, $P-Q$ is positive semi-definite: $P\succeq Q\Leftrightarrow P-Q\succeq 0$.
@@ 2.2 Definition of the Fisher-Rao distance and Fisher-Rao geodesics
Those Fisher information matrices $I(\lambda)$, $I(\lambda^{\prime})$ and $I(\lambda^{\prime\prime})$ define the same underlying Fisher Riemannian metric $g^{\mathrm{Fisher}}$ and length element. The Fisher information matrix for bivariate normal distributions was reported in [106]. See Appendix A for symbolic computing of the Fisher information matrices.
@@
Stigler [112] noticed that Hotelling historically first considered the Fisher-Rao distance in a handwritten note [60] dated in 1929. Hotelling considered the “space of statistical parameters” while Rao studied the “population space.” Interestingly, Mahalanobis mentioned the notion of “statistical field” in his celebrated paper introducing the Mahalanobis distance with tensor notations [31]. Nowadays, the term “statistical manifold” is used to either refer to a manifold of statistical models or to the pure dual geometric structure that was unravelled by Amari and Nagaoka [89, 6].
@@
Let $(\mathcal{X},\Sigma_{\mathcal{X}})$ and $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ be measurable spaces, and consider a stochastic kernel $K:\Sigma_{\mathcal{Y}}\times\mathcal{X}\rightarrow[0,1]$ [75]. A statistical distance $D(P_{1},P_{2})$ is said monotone [75] if $D(KP_{1},KP_{2})\leq D(P_{1},P_{2})$. The class of $f$-divergences [38] are monotone [75].
For example, consider the statistical model of categorical distributions on a sample space $\mathcal{X}$ of size $d$. Each categorical distribution is parameterized by a point $p$ on the $(d-1)$-dimensional standard simplex $\Delta_{d-1}$ sitting in $\mathbb{R}^{d}$. The Fisher-Rao distance between $p$ and $q$ [84] is
$\rho(p,q)=2\,\mathrm{arccosh}\left(\sum_{i=1}^{d}\sqrt{p_{i}q_{i}}\right),$
where $\mathrm{arccosh}(x)=\log(x+\sqrt{x^{2}-1})$ for $x\geq 1$. We may view categorical distributions $p=(p_{1},\ldots,p_{d})$ as normalized histograms with non-empty bins. Consider reducing the dimension $d$ of categorical distributions $p$ by merging some bins to get categorical distributions $p^{\prime}$ of $\Delta_{d^{\prime}-1}$ with $d^{\prime}<d$. This merging bin operation can be interpreted as a deterministic Markov kernel. Then we have $\rho(p^{\prime},q^{\prime})\leq\rho(p,q)$. In the extreme case of $d^{\prime}=1$ (i.e., merging all bins to a single bin of value $\sum p_{i}=1$), we have $\rho(p^{\prime},q^{\prime})=\rho(1,1)=0\leq\rho(p,q)$.
Property 1
The Fisher-Rao distances are monotone.
Proof:
We shall use the property of the Fisher information matrix. Let $Y=t(X)$ be a statistic. Then $I_{Y}(\theta)\preceq I_{X}(\theta)$ with equality if and only if $Y$ is sufficient [65]. Thus we have the following inequality for corresponding Fisher metrics: $g_{Y}(\theta)\preceq g_{X}(\theta)$, and the corresponding Fisher-Rao length elements $\mathrm{d}s_{Y}$ and $\mathrm{d}s_{X}$ are such that $\mathrm{d}s_{Y}\leq\mathrm{d}s_{X}$ because $\mathrm{d}s_{Y}^{2}={\mathrm{d}\theta}^{\top}g_{Y}(\theta){\mathrm{d}\theta}\leq\mathrm{d}s_{X}^{2}={\mathrm{d}\theta}^{\top}g_{X}(\theta){\mathrm{d}\theta}$ since $g_{X}(\theta)-g_{Y}(\theta)\succeq 0$ (positive semi-definite). Let $\gamma_{X}$ and $\gamma_{Y}$ be the geodesics with respect to Fisher metrics $g_{X}$ and $g_{Y}$, respectively. We have
$\rho_{Y}(\theta_{1},\theta_{2})=\int_{\gamma_{Y}}\mathrm{d}s_{Y}\leq\int_{\gamma_{X}}\mathrm{d}s_{Y}\leq\int_{\gamma_{X}}\mathrm{d}s_{X}=\rho_{X}(\theta_{1},\theta_{2}).$
Thus we have proved the monotonicity property of the Fisher-Rao distances: $\rho_{Y}(\theta_{1},\theta_{2})\leq\rho_{X}(\theta_{1},\theta_{2})$. $\square$
@@ 3.2 Fisher-Rao distances for products of stochastically independent models
For $m$ independent statistical models $\mathcal{M}_{1}=\{p_{\theta^{1}}(x_{1})\}$ , $\ldots$, $\mathcal{M}_{m}=\{p_{\theta^{m}}(x_{m})\}$ with corresponding variables $x_{i}\in\mathbb{R}^{d_{i}}$ and parameters $\theta^{i}\in\mathbb{R}^{m_{i}}$ (with $d=\sum_{i}d_{i}$ and $m=\sum_{i}m_{i}$), and Fisher-Rao’s distances $\rho_{\mathcal{M}_{1}}$, $\ldots$, $\rho_{\mathcal{M}_{m}}$, we have the Fisher-Rao distance on the product model =$\mathcal{M}=\{p_{\theta}(x)\propto\prod_{i=1}^{m}p_{\theta^{i}}(x_{i}){\ :\ }\theta\in\mathbb{R}^{m}\}$ which is given by
@@ 3.3 The Fisher-Manhattan upper bound for the Fisher-Rao distances
Figure 2: Illustration of the Fisher-Manhattan upper bound for a bidimensional statistical model ($m=2$): We calculate the 1D Fisher-Rao distances along the edges of the square and then find the shortest path on the weighted graph between node $\theta_{0}$ and node $\theta_{1}$.
@@
Figure 3: Illustrating the Fisher-Manhattan upper bound for the Fisher-Rao manifold of univariate normal distributions $\mathcal{N}$ ($m=2$): The Fisher-Rao geodesic between $N(\mu_{0},\sigma_{0})$ and $N(\mu_{1},\sigma_{1})$ is shown in red. The Fisher-Manhattan upper bound is derived from the straight-edge square $((\mu_{0},\sigma_{0}),(\mu_{0},\sigma_{1}),(\mu_{1},\sigma_{1}),(\mu_{1},\sigma_{0}))$. The Fisher-Manhattan upper bound is $\min(\rho((\mu_{0},\sigma_{0}),(\mu_{0},\sigma_{1}))+\rho((\mu_{0},\sigma_{1}),(\mu_{1},\sigma_{1})),(\mu_{0},\sigma_{0}),(\mu_{1},\sigma_{0})+\rho((\mu_{1},\sigma_{0}),(\mu_{1},\sigma_{1})))$, i.e., the minimum length of the two green and blue straight-edge path lengths. Tissot indicatrices of the Fisher metric are displayed at 2D grid locations. The submanifolds $\mathcal{N}_{\sigma_{0}}$ and $\mathcal{N}_{\sigma_{1}}$ are not totally geodesics and the submanifolds $\mathcal{N}_{\mu_{0}}$ and $\mathcal{N}_{\mu_{1}}$ are totally geodesics.
@@ 5 Tight upper bounds on Fisher-Rao distances: The case of Hessian metrics
Figure 6: Diagram illustrating the canonical relationships between metric tensors $g$ and induced connections: A metric-compatible Levi-Civita connection $\nabla^{g}$ is always induced by a metric $g$. When the metric is Hessian, we can further associate to torsion-free flat connections $\nabla$ and $\nabla^{*}$ such that $\nabla^{g}=\frac{\nabla+\nabla^{*}}{2}$.
A metric tensor $g$ induces a unique torsion-free metric connection called the Levi-Civita connection $\nabla^{g}$. When the metric $g$ is Hessian, there exists two torsion-free flat connections $\nabla$ and $\nabla^{*}$ such that $g=\nabla df=\nabla^{*}df^{*}$ for a pair of dual potential functions $f$ and $f^{*}$. Furthermore, we have $\frac{\nabla+\nabla^{*}}{2}=\nabla^{g}$, i.e., the metric is preserved by dual parallel transport [4]. Figure 6 illustrates the canonical relationships between the metric tensor $g$ and induced connections.
Dually flat spaces in information geometry [11, 4] are special Hessian manifolds which admit global coordinate charts and dual potential functions which are Legendre-type convex functions when expressed in the dual $\nabla$- and $\nabla^{*}$-affine coordinate systems. When the Fisher metric $g^{\mathrm{Fisher}}$ of a statistical model =$\mathcal{M}=\{p_{\theta}{\ :\ }\theta\in\Theta\}$ is a Hessian metric, the Fisher-Rao manifold $(\mathcal{M},g^{\mathrm{Fisher}})$ is called a Hessian manifold [108]. When a Riemannian metric is Hessian, Recall that the Fisher metric is defined in the global chart by the Fisher information matrix $I(\theta)=-E_{p_{\theta}}\left[\nabla^{2}_{\theta}\log p_{\theta}(x)\right]$. When the Fisher metric is Hessian, we have $G(\theta)=I(\theta)=\nabla^{2}F(\theta)$ for a strictly convex and smooth potential function $F(\theta)$. A dually flat space admits as a canonical divergence a Bregman divergence $B_{F}(\theta_{1}:\theta_{2})=F(\theta_{1})-F(\theta_{2})-(\theta_{1}-\theta_{2})^{\top}\nabla F(\theta_{2})$ induced by a potential Legendre-type convex function $F(\theta)$ (defined up to a modulo affine terms), and two torsion-free flat connections $\nabla$ and $\nabla^{*}$ such that $\frac{\nabla+\nabla^{*}}{2}$ coincides with the Levi-Civita connection. These dual connections $\nabla$ and $\nabla^{*}$ coupled to the Hessian Fisher metric $g^{\mathrm{Fisher}}$ allows to connect any two points $p_{\theta_{1}}$ and $p_{\theta_{2}}$ on the manifold either by $\nabla$-geodesics $\gamma_{p_{\theta_{1}}p_{\theta_{2}}}$ or by $\nabla^{*}$-geodesics $\gamma_{p_{\theta_{1}}p_{\theta_{2}}}^{*}$ (Figure 7).
@@
Many statistical models yield Fisher metrics: For example, exponential families [15], mixture families [4], all statistical models of order $m=2$ [80, 25, 8], and some elliptical distribution families [80, 81], etc. In general, there are obstructions for a $m$-dimensional Riemannian metric $g$ to be Hessian [5]. Thus it is remarkable that exponential and mixture families of arbitrary order yield Hessian Fisher metrics.
One can test whether a metric $g$ given in a coordinate system $\theta$ by $G(\theta)=[g_{ij}(\theta)]$ is Hessian or not as follows:
Property 2 (Hessian metric test [108])
A $m$-dimensional Riemannian metric $g$ expressed in the $\theta$-coordinate system is Hessian if and only if:
@@
Beware that this property holds for a considered parameterization $\theta$ of the metric $g$.
In general, when we change the coordinates of $\lambda$ to $\theta$, we consider the following covariant transformation rule of the metric tensor:
$g_{\theta}(\theta)=\mathrm{Jac}_{\theta}\lambda^{\top}\,g_{\lambda}(\lambda(\theta))\,\mathrm{Jac}_{\theta}\lambda,$ (20)
where $\mathrm{Jac}$ denote the Jacobian matrix: $\mathrm{Jac}_{\theta}\lambda=\left[\frac{\lambda_{i}}{\theta_{j}}\right]_{ij}$. Thus to check that a metric $g(\lambda)$ given in some parameterization $\lambda$ is Hessian or not, we need to find a parameterization $\theta(\lambda)$ such that $g_{\theta}(\theta)=\mathrm{Jac}_{\theta}\lambda^{\top}\,g_{\lambda}(\lambda(\theta))\,\mathrm{Jac}_{\theta}\lambda=\nabla^{2}F(\theta)$.
@@
Property 3 (1d Fisher metrics are Hessians)
@@
A 2D Hessian metric $g$ can always be expressed by a diagonal matrix in some coordinate system by a change of parameterization: Indeed, let $\theta=(\theta_{1},\theta_{2})$ and $\eta=(\eta_{1},\eta_{2})$ be the dual affine $\nabla$- and $\nabla^{*}$-coordinate systems of the flat connections $\nabla$ and $\nabla^{*}$. The Fisher metric is expressed in $\theta$ as $g(\theta)=\nabla^{2}F(\theta)$ and in $\eta$ as $g(\eta)=\nabla^{2}F^{*}(\eta)$. Then we can consider the mixed parameterizations $\xi=(\theta_{1},\eta_{2})$ or $\xi^{\prime}=(\eta_{1},\theta_{2})$ and obtain diagonal matrices for the Fisher information matrices with respect to those mixed parameterizations [4, 83].